ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 57S10 Repovš, Dušan, University of Ljubljana, Slovenia On Banach-Mazur compacta and their geometric properties We shall present a joint work with Sergei M. Ageev, the main result of which asserts that Q(n)-E is a Hilbert cube manifold, for every n ≥ 2, where E is the Euclidean point of the Banach-Mazur compactum Q(n). As a corollary we prove that Q(2) is not a homogeneous space.